OpenStudy (anonymous):

Carla is making ramps of different lengths. The ramps are made with rectangular and square wooden boards. The relationship between the length of the ramp in feet, f(n), and the number of square boards used, n, is represented by the function below. f(n) = 4n + 5 , where 20 ≤ n ≤ 30 Which statement is correct about the range of the length of the ramp for the given domain? It varies from 20 to 30. It varies from 20 to 125. It varies from 80 to 120. It varies from 85 to 125.

OpenStudy (anonymous):

plug in 20 and 30

OpenStudy (anonymous):

and solve

OpenStudy (anonymous):

to get range

OpenStudy (anonymous):

I think you just have to plug in 20 and 30 into the function f(n) = 4n + 5

OpenStudy (anonymous):

d. simply put lower and upper values of n in equation. . .

OpenStudy (anonymous):

@theyatin Your answer is invalid

OpenStudy (anonymous):

so B or which one

OpenStudy (anonymous):

romero u said same thing what i have said u are saying u are wrong or u need credit lol dnt mind

OpenStudy (anonymous):

??? which one

OpenStudy (anonymous):

f(n)=4(30)+5=125 and f(n)=4(20)+5=85 so the answer is d.

OpenStudy (anonymous):

@theyatin No one really cares if you can do the work. That is not the point. The point is trying to get the asker to do it and by simply giving out the answer you do nothing in helping the asker. Point the asker asked which answer it was from your explanation.

OpenStudy (anonymous):

easy romero bro easy i told her how to do it along with answer look if you didnt had. why people here are so fight seeking. . .

OpenStudy (anonymous):

There is no reason for you at all to give out the answer in any question. If you are you are doing a bad job of explaining. BTW that is against the code of conduct in OS I've seen you do it several times already

OpenStudy (anonymous):

easymy lord please i beg your pardon i will never ever will do this crime of telling answer along with explanation don't execute me for this.... please............... ;'(

OpenStudy (anonymous):

what kind of nerdy sht is this?